OpenStudy (anonymous):

Please show that the space of bounded functions on a set X, l^∞(X)={f∈R^X |There is M∈R so that for all x∈X, we have |f(x)|≤M} is a subspace of R^x. Or give me a hint on this problem? Thanks

OpenStudy (anonymous):

So, prove that: \[ I^\infty(X)=\{f\in\mathbb{R}^X|\exists M\in\mathbb{R} \text{ s.t. }\forall x\in X, |f(x)|\le M\} \]Is a subspace of \(\mathbb{R}^x\)? (Just to make sure I'm understanding the problem, correctly. And, what do you mean when you say \(f\in \mathbb{R}^X\)? \(X\) is a set, do you mean its cardinality?

OpenStudy (anonymous):

I'll try to rewrite the question... "Show that on a set X, the space of bounded functions \[l^∞(X) = \left\{ f \in \mathbb{R}^X | \exists M \in \mathbb{R}; s.t. \forall x \in X, |f(x)|≤M \right\},\] is a subspace of \[\mathbb{R}^X.\] In terms of what I mean when I say: \[f \in \mathbb{R}^X\] ... well that's just what it says in this homework question and the whole question is confusing me, so I'm not going to be of very much help, unfortunately. Sorry!

OpenStudy (anonymous):

Screen shot of the question...

OpenStudy (anonymous):

Hmm... I think I get what your professor means by \(f\in\mathbb{R}^X\). I think it means: \[ f:X\mapsto \mathbb{R} \]It's a little odd he or she didn't define that in class or on the homework... All right, so, we know that: \[ l^\infty(X)\subset\mathbb{R}^X \]By definition. Now, let's declare two functions \(f, g \in l^\infty(X)\). We know that, for such to be a subspace of \(\mathbb{R^X}\), \(f+g\in l^\infty(X)\) and, \(\forall c\in\mathbb{R}\), \(cf\in l^\infty(X)\). Let's prove these statements: If: \[ f, g \in l^\infty(X)\implies\\ \exists M_f\in \mathbb{R}, s.t., \forall x\in X, f(x)\le M \]We declare the same for \(M_g\), for function \(g\). Then: \[ f+g\le M_f+M_g \]Since: \[ M_f+M_g\in\mathbb{R} \wedge f+g\in R^X\implies f+g\in l^\infty(X) \]And, we do the same, for some \(f\in l^\infty(X)\), \(c\in\mathbb{R}\): \[ \forall x, |f(x)|\le M\implies\\ c|f(x)|\le cM \]And, since: \[ c\cdot f(x)\in\mathbb{R}^X \]We are done, thus, \(l^\infty(X)\) is a subspace of \(\mathbb{R}^X\)

OpenStudy (anonymous):

Sorry, I forgot to put absolute values around some things, but I'm sure you get what I mean.

OpenStudy (anonymous):

Oh, and, for the last part: \[ cM\in\mathbb{R} \]